# Story Generation from Visual Inputs: Techniques, Related Tasks, and Challenges

### v1

@@ 7. Conclusions
This survey reviewed the field of story generation, explaining its development and current methods. It started by explaining basic story elements and early techniques, which mostly involved manually created rules. We looked at the datasets and evaluation metrics used to judge the quality of the stories. The discussion pointed out the need for better metrics to assess complex tasks like story generation accurately. The document also covered related areas like image and video captioning and VQA, discussing how improvements in these fields have helped enhance VSG.
The survey also looked at the latest developments in VSG and discussed various models that use deep learning. It started with basic models using simple encoder-decoder setups and went on to more advanced systems that use transformers. Each model was described with a focus on how they handle issues like making the story coherent, keeping it relevant to the context, and combining detailed visual and textual information.
@@
7.1. Exploring Large Language Models
@@
The field of VSG using LLMs is under-explored, with only a few works delving into this domain. To the best of our knowledge the existing works predominantly rely on GPT-2, which is an outdated LLM when compared to models such as Mixtral (team, 2023) which is the best performing open-source model on MT-Bench or GPT-4 which is the best performing proprietary model. The scarcity of the integration of LLMs with visual elements presents an opportunity for further research, exploring the potential of these models in generating more coherent and contextually relevant stories.
7.2. Automatic Evaluation Metrics
@@
The need for robust and efficient automatic evaluation metrics is emphasized in our survey. While existing metrics like METEOR and CIDEr have been widely adopted, they do not fully capture the nuances of story quality. The metrics mostly focus on the similarity between the generated text and the reference text. A story can be retold in various ways, employing different perspectives while preserving the original message. Consequently, a generated text might deviate significantly from the reference text in terms of specific wording and still deliver the same story in a coherent way. Thus, while these metrics provide valuable insights into certain dimensions of text generation, they did not fully encompass the multifaceted nature of storytelling. As LLMs continue to improve, they show potential in replacing human annotators in many tasks (Gilardi et al., 2023) and so metrics to evaluate stories could be developed based on LLMs. We advance that these metrics could be based on the existing systematic scoring systems (Zheng et al., 2023) that were developed to evaluate the performance of LLM-based chat bots.
7.3. Decomposition of Story Generation
@@
Addressing the complexity of stories, we underscore the importance of decomposing the story generation process into modular components. Each module can focus on specific attributes such as author goals, character interactions, consistency, etc. This decomposition has already been explored with modeling characters in CharGrid (Yu et al., 2021) and the modeling of characters and setting in HEGR (Zheng et al., 2021). We believe that explicit modeling of other story elements and abstract concepts such as emotions and culture, possibly using hypergraphs or similar structures could further enhance the quality of generated stories.
7.4. Hybrid Systems and Knowledge Integration
@@
Building on the decomposition concept, hybrid systems, combining planning-based techniques and machine learning approaches, present an intriguing direction. Hyper-graphs, as seen in HEGR (Zheng et al., 2021), could possibly be a starting point as a deep learning approach to model the planning process replacing the traditional handcrafted planning systems. The integration of multiple knowledge sources can enrich the diversity and quality of generated stories, paving the way for more robust and context-aware narrative creation.
@@ Acknowledgements.
Daniel A. P. Oliveira is supported by a scholarship granted by Fundação para a Ciência e Tecnologia (FCT), with reference 2021.06750.BD. Additionally, this work was supported by Portuguese national funds through FCT, with reference UIDB/50021/2020.

### v2

@@ 7 Conclusions
This survey reviewed the field of VSG, explaining its development and current methods. It started by explaining basic story elements, datasets and evaluation metrics used to judge the quality of the stories. The discussion pointed out the need for better metrics to assess complex tasks like VSG accurately. The document also covered related areas like image and video captioning and VQA, discussing how improvements in these fields have helped enhance VSG. The discussion then moved on to the evolution of VSG methods, starting with early approaches that used CNNs and RNNs to methods based on LLMs and hypergraphs. Additionally, we examined emerging real-world applications that demonstrate the practical potential of these technologies in domains such as automotive systems and location-aware storytelling.
@@
7.1 Exploring Large Language Models
@@
The field of VSG using LLMs is under-explored, with only a few works delving into this domain. To the best of our knowledge the existing works predominantly rely on GPT-2, which is an outdated LLM when compared to models such as Qwen3-235B-A22B Yang et al. (2025) which is the best performing open-source model on Chatbot Arena or Gemini 2.5 Pro Comanici et al. (2025) which is the best performing proprietary model. The scarcity of the integration of LLMs with visual elements presents an opportunity for further research, exploring the potential of these models in generating more coherent and contextually relevant stories.
7.2 Automatic Evaluation Metrics
@@
The need for robust and efficient automatic evaluation metrics is emphasized in our survey. While existing metrics like METEOR and CIDEr have been widely adopted, they do not fully capture the nuances of story quality. The metrics mostly focus on the similarity between the generated text and the reference text. A story can be retold in various ways, employing different perspectives while preserving the original message. Consequently, a generated text might deviate significantly from the reference text in terms of specific wording and still deliver the same story in a coherent way. Thus, while these metrics provide valuable insights into certain dimensions of text generation, they did not fully encompass the multifaceted nature of storytelling. As LLMs continue to improve, they show potential in replacing human annotators in many tasks Gilardi et al. (2023) and so metrics to evaluate stories could be developed based on LLMs. We advance that these metrics could be based on the existing systematic scoring systems Zheng et al. (2023) that were developed to evaluate the performance of LLM-based chat bots.
7.3 Decomposition of Story Generation
@@
Addressing the complexity of stories, we underscore the importance of decomposing the story generation process into modular components. Each module can focus on specific attributes such as author goals, character interactions, consistency, etc. This decomposition has already been explored with modeling characters in CharGrid Yu et al. (2021) and the modeling of characters and setting in HEGR Zheng et al. (2021). We believe that explicit modeling of other story elements and abstract concepts such as emotions and culture, possibly using hypergraphs or similar structures could further enhance the quality of generated stories.
7.4 Hybrid Systems and Knowledge Integration
@@
Building on the decomposition concept, hybrid systems, combining planning-based techniques and machine learning approaches, present an intriguing direction. Hyper-graphs, as seen in HEGR Zheng et al. (2021), could possibly be a starting point as a deep learning approach to model the planning process replacing the traditional handcrafted planning systems. The integration of multiple knowledge sources can enrich the diversity and quality of generated stories, paving the way for more robust and context-aware narrative creation.
7.5 Real-World Deployment and Context Integration
The transition from academic research to practical applications represents a crucial next step for VSG technologies. Recent work in location aware storytelling and automotive applications demonstrates the potential for deploying these systems in real-world scenarios. Future research should focus on developing robust systems that can integrate multiple contextual information sources including location, temporal data, and real-time sensor information while maintaining computational efficiency for deployment in resource constrained environments such as mobile devices and embedded systems. The success of these real-world applications will ultimately determine the broader impact of VSG research on society.
@@ Acknowledgements.
[…]